solve the following equations by cramer's rule: x+y+z=9, 2x+5y+7z=52, 2x+y-z=0
Answer by ssasr(5)

x+y+z=9
2x+5y+7z=52
2x+y-z=0
delta= 1 1 1
2 5 7
2 1 -1
=-4
delta x=
9 1 1
52 5 7
0 1 -1
=-4/-4=1
delta y=
1 9 1
5 52 7
1 0 -1
=-12/-4=3
delta z=
1 1 9
7 5 52
-1 1 0
=-20/-4=5
ANSWER=x=1,y=3,z=5

_____________M.Srikanth
XI std